2ćŘěŊcďŘÃêŘôÂœ ø ^{1}ěŚŞ ÀîcďŘÃêŘôÂœ

(79174904) ÀśúÔĺŘ ćŘÀúŞ ďŘÃêŘôÂœ ücĺÀĂúõ ğc ďć ýŘÂŞ ĆŐŘďŘ ťcúŻ

öŚþÂČõ Âţîć :ćŚţĺŘ 1382 ÂČŮ 17

Open Source Software^{1} Free Software^{2}

ùÀČØİ

ďć ÀþŚł Ćî ,ÀĂłŚŞüõ ďŘÃêŘôÂœ ücĺÀĂúõ űc ¡Ş ďć ÀþÀŻ ŚţśÆœ ôcĄúÔõ øć ćŘěŊcďŘÃêŘôÂœ ø ěŚŞ ÀîcďŘÃêŘôÂœ ýùÀœďŘćùŚÚœ ýŚûùÀþŘ ĆŞ ÂÊţ£õ ĆŞ üûŚÚœ ěŘ ÅŢ .ÀĂłŚśœ ű¡śõ SSþŘ ŚŞ ÍśŮÂõ öŘÀĂİ ñøŘ ù cŚÚœ Ć î ü ŮøŚ Ô ţ õ ýcďŚ ź Ů ñcÀ õ ěŘ ,À œďøŊü õ ćĄ Żø Ć Ş ďŘÃ êŘôÂ œ À cČ óĄ Ů À cœøď Ûcî ďć Ć î ü óĄ ¡ Ů ø Ś ûöŊ SSþŘ ďć ďŘÃêŘôÂœ ýĆãĺĄŮ ýŘÂŞ Ćî ćďŘć öŊ ěŘ öŚÈœ ĆÞû ,öŚłŰøŚÔţõ öcŚÆþĄœĆõŚœÂŞ ŚŮ ,ÀĂśÜÏüõ üûŚÚœ ěŘ ÅŢ ĆţłĄœ SSþŘ ďć .ÀłŚŞüõ ŚûùÀþÀŢ SSþŘ ŚŞ ŚĂłŊ ďcŘÃêŘôÂœ ücĺÀĂúõ Ý cÜä ĆŞ ěŚČœ ŚûñÀõ Ûc§ŘÂõ ěŘ Œþ Âû ýŘÂŻŘ ücÚœĄÚİ ÝČĂîüõ üãĺ ,ôĄúÔõ øć SSþŘ ýŚûŰøŚÔŮ ø ÓþďŚãŮ ĆŞ üóŚÞŻŘ ýŚŻĆŞ ćÂî ÝČûŘĄą üãĺ řČŮÂŮ SSþŘ ĆŞ .ÝČĂî üêÂãõ ěŚŞ Àî ýcŚČœć ďć Řď ďŘÃêŘôÂœ ücĺÀĂúõ À cœøď SSþŘ ýcěŚĺùćŚČŢ ĆŞ ,üãìŘø ýŚûùęøÂŢ ěŘ üþŚûñŚůõ öcŚČŞ ŚŞ ,ÝČĂî ćďĄąÂŞ ùÀþÀŢ SSþŘ ŚŞ ýďĄßŮ ďć ĆîöŊ

.ÝþćÂð ŒþćÃœ Śûľøď

ĆõÀÖõ 1

ěŘ ÂČè üÜõŚä Ý cţÆČĺ ŚŞ Ćî ÅîÂû .ÀĂţÆČœ ŚĂłŊŚœ ôcĄúÔõ øć ÂÚþć ćŘěŊcďŘÃêŘôÂœ ø ěŚŞ ÀîcďŘÃêŘôÂœ .ćďŘć áĚÏŘ ŚûďŘÃêŘôÂœ SSþŘ ťc ČÞûŘ ø ćĄŻø ěŘ Œł öcøÀŞ ÀłŚŞ ĆţłŘć ďŚîøÂĺ ^{1}ěøÀĂþø ťêŚĺøÂØþŚõ ÂŞ ýďøÂõ ĆØÜŞ ,ćÂî ÝČûŘĄ£œ ùďŚłŘ ŚûöŊ ýcÂČðÛØł ýùĄ¡œ ø ĆČóøŘ Óc þďŚãŮ ĆŞ ĆţłĄœ SSþŘ ďć ŘÁó ďcŘÃêŘôÂœ À cČóĄŮ À cœøď ø ĆţÈðÂŞ ñŚ§ öcŚõě ĆŞ öŊ ěŘ ÅŢ .ťłŘć ÝČûŘĄą öŚłÀþŘĄê ø ŚûöŊ ćcĄŻø ťc Üä ďŘÃêŘôÂœ ücĺÀĂúõ À cœøď ĆŞ ýÂÊţ£õ ù cŚÚœ öŊ ěŘ ÅŢ .ćÂî ÝČûŘĄą üĺďÂŞ Řď ěøÂõŘ ýĆãõŚŻ ďć ćŘěŊ ŚûöŊ ýcÀĂŞĆţĺć ø ÕêĄõ ýùęøÂŢ ÀĂİ ĆŞ üóŚÞŻŘ üûŚÚœ Åşĺ .ćÂî ÝČûŘĄą ěŚŞ Àî ýŚûďŘÃêŘôÂœ ďć ŚúĂŮĆœ ďŘÃêŘôÂœ ñcÀõ SSþŘ ,ŰďŚźŮ Â cÑœĆÎÖœ ěŘ Ćî ÝČûć öŚÈœ ćÂî ÝČûŘĄą üãĺ ďŚî SSþŘ ŚŞ .ÝČĂîüõ

.ťêÂð ÀûŘĄą üÈČŢ ŚûöŊ ěŘ ďŚČÆŞ ,ćĄł ĆţêÂð ďŚî ĆŞ ťĺďć ÂðŘ ĆØÜŞ ,ćďŘÀœ ÂÚþć ácŘĄœŘ ÂŞ üÔãő

2ěŚŞ Àî ýùÀûŚãõ 2

ýŚûďŘÃêŘôÂœ ÂŮĄČşõŚî ýcćŚä Â cŞďŚî Âû ,ćĄłüõ ťþŚäď ŚûöŊ ďć üşî Õ c§ öcĄœŚì Ćî öďÀõ ýŚûďĄÈî ďć Řď ľÂŮĄČşõŚî ÂŞďŚî ťĺŘ SSØÞõ .ÀĂîüÞœ ùćŚÔţĺŘ ÂÚþć ø ťĺŘ ùÀþÂą Ûśì ŚûñŚĺ Ćî ćďŘć üœŘøŘÂê ŚûďŘÃêŘôÂœ .ćÂî ÀĂûŘĄ£œ ďŚî ÂÚþć ÇþŚûĆõŚœÂŞ ø ,ÀĂî ùćŚÔţĺŘ ýÂÚþć ñcÀõ ěŘ Śþ ÀłŚŞ ùćŘć ŚÖŮďŘ SSØÞõ .ÀûÀœ ôŚźœŘ ÀûŘĄąüõ ÂŞďŚî Ćî Řď ýďŚî ĆõŚœÂŞ ťĺŘ SSØÞõ .ÀĂłŚŞ ùÀł üÞþÀì ťĺŘ SSØÞõ ñĄŢ ďŘÃêŘôÂœ öŚÞû ěŘ ýÂÚþć ýĆ£Æœ ťc ŞŚŞ ÀûŘĄ£œ ø ÀłŚŞ ùÀþÂą ÂÚþć Â cŮĄČşõŚî ÀĂİ Śþ øć ťĺŘ Řď ľďŚî ěøÂõŘ ,ťĺŘ ùÀł ùćŘć ñĄŢ ľť dŞŚŞ ÇČŢ ŚûñŚĺ Ćî ýďŘÃêŘôÂœ ,ÀłŚŞ Ćî ÛČóć Âû ĆŞ .ćěŘćÂşŞ

?Àûć ăď ëŚÔŮŘ SSþŘ ťĺŘ ôěĎ ŚþŊ .ÀûćüÞœ ôŚźœŘ

Ćİ ?ÀČûć ŚÖŮďŘ üœŚźõ Řď ćĄą ďcŘÃêŘôÂœ ťĺŘ ôěĎ ùŚðÂû ŚŮ ÀČţłŘć Řď Õ§ SSþŘ ÂðŘ Àłüõ Ćİ ?ÀþćŘćüõ ÂČČçŮ üœŚźõ ÃČœ Řď ďŘÃêŘôÂœ ýĆ£Æœ ÀþćŘć ÝţÆČĺ Â cČČçŮ Mac ĆŞ PC ěŘ üţìø ÂðŘ Àłüõ Çţĺďć ÀČûÀŞ ÀČţÆœŘĄŮüõ ,ćĄśœ ýĄì ôěĎ ýùěŘÀœŘ ĆŞ Śþ ćÂîüÞœ ďŚî ďŘÃêŘôÂœ üţìø ÂðŘ Àłüõ Ćİ ľŘùÀœěŚĺ ťc îÂł ĆîöŊ ěŘ ÅŢ öŚĂİÝû ďŘÃêŘôÂœ ÂðŘ Àłüõ Ćİ ?ÀČĂî Çţĺďć öŚŮćĄą üţ§ Śþ ,ÀĂĂî ýŚŻ ĆŞ Řď öŚŮďŘÃêŘôÂœ ÀČţÆœŘĄŮüõ ÂðŘ Àłüõ Ćİ ?Àłüõ üœŚśČţÈŢ Ýû ěŚŞ ,Àł ŽďŚą ŰďŚźŮ ďcŘěŚŞ ěŘ ŰďĄň SSþŘ ďć ?ÀČĂî řÊœ öŚŮÛÞ§ÛcŞŚì Â cŮĄČşõŚî ø ,ñÃĂõ ,ďŚî Ûc¡õ Â cŮĄČşõŚî ýøď ,ÂŮĄşõŚî Œþ ,ÀœďŘć ćĄŻø üìĄÖ§ ÅŢ .ÀþćÂîüõ ùćŚÔţĺŘ ÀþćĄŞ ùÀþÂą ÇČŢ ŚûñŚĺ Ćî ýďŘÃêŘôÂœ ěŘ ěĄĂû ĎŚÞţ§Ř

.Àûćüõ ďŘÂì öŚŮďŚČţąŘ ďć ěŚŞ Àî Ćî ÓþÂãŮ Řď üÊ£Èõ ëcĄÖ§ Ćî ,ťĺŘ ÂŮĄČşõŚî Â cŞďŚî ýŘÂŞ üìĄÖ§ ýĆäĄÞźõ ěŚŞ Àî ýùÀûŚãõ Ćî öŚœŊ .ćĄł ĆţąŚĂł ěŚŞ Àî öcŘĄĂä ĆŞ ŚŮ Àûć ďŘÂì öŚŮďŚČţąŘ ďć ÀþŚŞ ďŘÃêŘôÂœ ^{3} ěcĄźõ Ćî ÀĂîüõ ëcĄÖ§ ĆŞ öŘÂŞďŚî ø ÀĂŞŚþüõ ÛØÈõ Řď ÀœěŚĺüõ Ćî öŚœŊ ŚŞ ťŞŚìď ÀœěŚĺüÞœ ěŚŞ Àî Řď ćĄą ýŚûĆõŚœÂŞ ø ^{4}ÅØĂ cČó ÛcõŚä Ý cţÆČĺ À cĂœŚõ üþŚûĆõŚœÂŞ .ÀœŘĆţłŘćüõ ÀþŚŞ ùďŘĄÞû Ćî ÀĂŞŚþüõ ťĺć ýÀþÀŻ ýÂŮćøÀ¡õ ěcĄźõ Ćî ÀœŘùÀł ýÂÚþć ýŚûďŘÃêŘôÂœ SS cþÃÚþŚŻ ø ÀœŘĆţÈð ÿþŘď ďŚČÆŞ ^{5}šČØcĺťœ Â cðďøÂõ ćĄĺ ŚûöŊ ľc ÂţÆð ýĎŚŞ ťc äÂĺ ěŘ ÀĂĂîüõ ùćŚÔţĺŘ ěŚŞ Àî ýŚûďŘÃêŘôÂœ ěŘ Ćî üþŚûťîÂł .ÀœďŘć ěcŚČœ âcêď ýŘÂŞ Ćî ýćŘÂêŘ Íc ĺĄŮ öŊ ěŘ ýďŚČÆŞ ø ,ťîÂł ÀĂİ ýcďŚØÞû Íc ĺĄŮ ĎĄÞãõ Ćî ,ÀœÂŞüõ

Microsoft Windows^{1} The Open Source Definition^{2}

.ćĄłüõ ôŚźœŘ ÀĂĂîüõ Řď ďŚî SSþŘ öŚłüÊ£ł

,ÀĂûćüõ ĆõŘćŘ ćĄą ďcŚî ĆŞ ÀœÀþÂêŊ Řď ÅØĂČó öĄİ üŮĎĄÊ¡õ Ćî üœŚśÜÏøŘć Ćî öŚĂİÝû ěŚŞ Àî Ćî üìĄÖ§ ćcĄŻø ŚŞ ŚúĂŮ SSþŘ ø ,ÀœěŘćÂŢüõ öŚłćĄą ÛcČõ ĆŞ ďŘÃêŘôÂœ ýĆãĺĄŮ ÃČœ ŚûťîÂł ŚûťäŚĺ ñcĄÊ¡õ ĆîSSþŘ ěŘ ÂŮĄČşõŚî öcŚÆþĄœĆõŚœÂŞ ÍcĺĄţõ Â cÈì .ťĺŘ ĆţÈð ÕÖ¡õ ćďøŊüõ ÝûŘÂê ,ÀłŚŞ ùÀł öŚłÀþŚä ýÃČİ ÇþŘěŘ ďć ĆîöŊüŞ ÀĺÂŞ ľøÂê ĆŞ ďŘÃêŘôÂœ řc §Śň Íc ĺĄŮ öŚłľĚŮ ğc ŚÆ§Ř ěŚŞ Àî ýŚûďŘÃêŘôÂœ ĆŞ ŒÞî ďć öŚÆþĄœĆõŚœÂŞ SSČÞû .ÀĂĂîüõ ŰďŚÖ§ ø ťìŚÞ§ ğc ŚÆ§Ř

:ÀœŘĆţêŚþ ťĺć ëĄÖ§ SSþŘ ĆŞ ÇþŘěŘ ďć ŘÂþě ÀĂĂîüõ üţ§Řď

.Śûüşî öŊ ďcŚÈţœŘ ø üşî Õ c§ •

.ÀłŚŞüõ öŊ ďć ÂČČçŮ ěcŚČœÇČŢ Ćî ,ďŘÃêŘôÂœ âcśĂõ À cî ĆŞ üĺďťĺć Õ c§ •

.ĆõŚœÂŞ đcĚňŘ ø ľÂţÆð Õ c§ •

ýĆ Þ û ŘÂ þě ,À Ă łŚ Şü õ Ý ú õ À Ă îü õ Œ Þ î ďŘÃ êŘôÂ œ ľc Â ţ Æ ð Ć Ş Ć î ü Æ î ýŘÂ Ş ëĄ Ö § SS þŘ Œþ ÀœŘĄŮüõ ÀûŘĄ£Ş Ćî ÅîÂû .ÀœďŘćüõ ùŚÚœ ĳÎĺ Œþ ďć Ýû ĆŞ ťśÆœ Řď öŚðÀĂûćľÂţÆð ýŚûďŘěŚŞ ĆŞ üŞŚþťĺć ýŘÂŞ ľÂţÆð ø ćĄŞ ÀĂûŘĄą SSČþŚŢ ŚûťÞČì ĆźČţœ ďć ,ÀłøÂÔŞ Řď ěŚŞ Àî ďcŘÃêŘôÂœ ťìø ěŚŞ Àî ďcŘÃêŘôÂœ Œþ ýùďŚŞďć áĚÏŘ řc Æî ýŘÂŞ Ćî ÅîÂû .Àł ÀûŘĄą ôŚźœŘ ÂŮâþÂĺ ÀþÀŻ øŘ üc œŚśČţÈŢ ěŘ Ćî ÀûÀŞ Řď ŠŚ£ţœŘ SSþŘ öŘÂŞďŚî ĆŞ ø ,ÀĂî üœŚśČţÈŢ Řď öŊ ÀœŘĄŮüõ ÀĂî éÂň ø ÀþÀŻ ýďŘěŚŞ ĆŞ üŞŚþťĺć ýŘÂŞ Řď ěŚŞ Àî ýĆõŚœÂŞ Œþ ÀœŘĄŮüõ üÆþĄœĆõŚœÂŞ Âû .ÀœćÂð ÀĂõùÂúŞ ýŘĆĂþÃû ťc ąŘćÂŢ ĆŞ ďĄśźõ ÀĂûćüõ ôŚźœŘ Řď ŚûďŚî SSþŘ Ćî üœŚÆî .Àûć ÂČČçŮ ÀþÀŻ ýŚûýÂţÈõ

.ÀĂţÆČœ öŚłďŚî ôcŚźœŘ ýùěŚŻŘ ýŘÂŞ

öŚúŻ Â cĺŘÂĺ ďć ťÆČœĄÞî ťc ÆØł Ćî üóŚ§ ďć ,ÀĂœŚõťÆČœĄÞî ýcÄŮŘÂţĺŘ SSþŘ ťc ČÖêĄõ ťc Üä .ťĺŘ ŰøŚÔţõ ÂÚþć ŰcĎĄÊ¡õ ćcŚÊţìŘ ŚŞ ĎĄňŘ ŰŚäĚÏŘ ćcŚÊţìŘ Ćî ťĺŘ SSþŘ ,ťĺŘ ćĄúÈõ ø ëÂŞ ýĆĂþÃû .ťĺŘ ÃČİŚœ ďŚČÆŞ ÂŮĄČşõŚî ýĆõŚœÂŞ À cĂœŚõ ,ŰŚäĚÏŘ ýŘĆØŮ öcćÂîüşî ýĆĂþÃû öcøÀŞ Řď öŚœ {ń \atop c} Âì Œþ öŘĄŮüÞœ ,ÛŞŚÖõ ďć .ťĺŘ ÂÑœéÂň ÛcŞŚì ĚõŚî öŊ ŰcŘÃČúźŮ íc ĚúţĺŘ

License^{3} Linux^{4} Netscape^{5}

.ćÂî üşî ćďŊ ďÀìöŚÞû SS cţłŘć

Ćż£þďŚŮ 3

üŮŚÖČÖ¡Ů ýďŘÃŞŘ ťêŚþ ùŘď ŚûùŚðÇœŘć ĆŞ ÂŮĄČşõŚî Ćî ŘÀţŞŘ .ťĺŘ üÞþÀì ďŚČÆŞ ćŘěŊcďŘÃêŘôÂœ ôcĄúÔõ Ćœ ,ÀĂţêÂðüõ ëĄÖ§ üÆþĄœĆõŚœÂŞ ÛcÞä ťc ŞŚŞ ŚûÅþĄœĆõŚœÂŞ ø ćĄŞ ğďťĺć ďć ĆœŘćŘěŊ ďŘÃêŘôÂœ .ćĄŞ ĆŞ áøÂł ŚûÅþĄœĆõŚœÂŞ Ćî ćĄŞ ťêŚþ ùŘď ŰďŚźŮ ýŚČœć ĆŞ ÂŮĄČşõŚî Ćî ŚûÀãŞ ŚúĂŮ .ŚûĆõŚœÂŞ ćcĄą ýŘÂŞ ücşî Âû ýŘÂŞ ýŘĆĂþÃû ťc êŚþďć ø ,öŚłďŘÃêŘôÂœ ĆŞ öŘÂŞďŚî ëcĄÖ§ öcćÂîćøÀ¡õ ÍcĺĄŮ ćĄą ěŘ ťþŚÞ§ ěŚèŊ 1984 ñcŚĺ ěŘ ø ^{6}SSÞóŚţĺŘ ćďŚżþď ÍĺĄŮ üĺŚČĺ üûŚðÀþć öcŘĄĂä ĆŞ ćŘěŊcďŘÃêŘôÂœ .ÀœćÂî öŊ SSÞóŚţĺŘ .ćÂî ěŚèŊ Řď ^{8}ĄĂð ýùęøÂŢ ø ćŘć ÛČØÈŮ Řď ^{7}ćŘěŊcďŘÃêŘôÂœ ćcŚČĂŞ ñŚĺ SSþŘ ďć SSÞóŚţĺŘ .Àł øŘ .ÀœÂśŞ ùÂúŞ öŚłýćŘěŊ ěŘ ÀþŚŞ ø ,ÀĂłŚŞ ĆţłŘć ýÂţÈČŞ ýcćŘěŊ ÀþŚŞ ôćÂõ Ćî ćŚúœ SSþŘ ÂŞ Řď ŋÂê ŚûöŊ ø ,ÀĂłŚŞ ĆţłŘć ÀþŚŞ öŘÂŞďŚî ýĆÞû ćĄŞ ÀÖţãõ Ćî ćÂî ü§ŘÂÏ Řď üìĄÖ§ ěŘ ýŘĆäĄÞźõ Åşĺ ĆŞ üłěďŘŚŞ ýŚûďŚî ćĄą SSÞóŚţĺŘ .ćÂî řŮÂõ GPL öŚÞû Śþ GNU General Public License ďć Řď öŊ ěŘ ^{10} ÅØÞþŘ ĄĂð À cĂõöŘĄŮ ďŚČÆŞ Â cðÇþŘÂþø ø ,^{9}ĄĂð C Â cÜþŚşõŚî Ćî ćŘć ôŚźœŘ ćŘěŊcďŘÃêŘôÂœ öcŘĄĂä ćŘěŊcďŘÃêŘôÂœ GPL ťc ¡Ů ÃČœ öŚÈþŘ Àł űäŚŞ ø ùćŚúœ ÂČŹŚŮ ýÂÚþć ďcŚČÆŞ ćcŘÂêŘ ďć ďŚŹŊ SSþŘ .ÀœŘĆţĺć ýďŚČÆŞ ÛcõŚł ěŚŞ Àî ýùÀûŚãõ Œł öcøÀŞ üóø ,ťĺŘ ùÀÈœ ùďŚłŘ ŚÞČÖţÆõ ŚŻžČû ĆİÂðŘ .ÀĂĂî ÀČóĄŮ

.ťÆœŘć øŘ ýŚûďŚî ěŘ üÖţÈõ Řď öŊ öŘĄŮüõ âìŘø ďć ø ťĺŘ SSÞóŚţĺŘ ýŚûùÀþŘ ěŘ

.ÀõŊ ćĄŻøĆŞ ^{11} SSČdŞć ôcŚœ ĆŞ ÅØĂČó ěŘ üãþěĄŮ ťc ĺŚČĺ âcŻÂõ öcŘĄĂä ĆŞ ŘÀţŞŘ ěŚŞ Àî ýùÀûŚãõ ěŘ ĚõŚî ÝţÆČĺ SSþŘ .ťĺŘ ÿþŘď ďŚČÆŞ ÃČœ ùěøÂõŘ Ćî ťĺŘ ÅØĂČó ýŚûÝţÆČĺ SSČóøŘ ěŘ üØþ SSČdŞć öŊ ďć Ćî ďŘÃêŘôÂœ Óc Üţ£õ ýŚûěĄźõ ácĄĂŮ ĆŞ ĆŻĄŮ ŚŞ ,ćĄŻø SSþŘ ŚŞ ,ćĄŞ ùÀł ĆţąŚĺ ćŘěŊ ýŚûďŘÃêŘôÂœ âcêď ýŘÂŞ .ćĄŞ ĆŻŘĄõ ÛØÈõ ŚŞ ťĺŘ ćŘěŊ ýďŘÃêŘôÂœ Ćİ ĆîSSþŘ Óc þÂãŮ ďć SSČdŞć ,Àłüõ ťêŚþ öŚõě ŚŞ Àłüõ üţ§Řď ĆŞ Ćî ÀœćÂî âőø Řď ýÀäŘĄì ^{12} ÅœÂ cŢ cğøÂŞeĆÜÞŻ ěŘ ,SSČdŞć ýŚÌäŘ ,ŚûôŚúŞŘ SSþŘ ŚûÀãŞ ÀäŘĄì SSþŘ .Ćœ Śþ ťĺŘ ćŘěŊ ďŘÃêŘôÂœ Ćî ťêÂð ÝČÞÊŮ ÀäŘĄì SSþŘ ŚŞ ďŘÃêŘôÂœ ěcĄźõ ýĆÆþŚÖõ öcŘĄĂä ĆŞ ,ćĄŞ ùÀł SSČdŞć ĆŞ Ćî üŮŘďŚłŘ écÁ§ ŚŞ ,^{13} ÀœĄÞþďcŒþďŘ cöcŘøŘÂê Œc Þî ŚŞ 1998 ñcŚĺ ďć

Richard Stallman^{6} Free Software Foundation^{7}

GNU^{8} GNU C Compiler (GCC)^{9}

GNU Emacs^{10} Debian GNU/Linux Distribution^{11}

Bruce Perens^{12}

.ťêŚþ ďŚÈţœŘ ěŚŞ Àî ýùÀûŚãõ

ćŘěŊcďŘÃêŘôÂœ ø (Open Source) ěŚŞ Àî ŰcďŚśä øć ýĆÎŞŘď ÀþŚÞœüõ ÙĂð öŚČõ SSþŘ ďć ĆİöŊ ücœŚãõ ĆŞ üÆČÜÚœŘ öcŚŞě ďć Open ýĆÞÜî Ćî ÀœćĄŞ ÀÖţãõ ýďŚČÆŞ ĆİÂðŘ .ťĺŘ (Free Software) ćcĄŻø ŚŞ Řď öŊ ÂÚþć ýùÀä ŚõŘ ,ććÂð ùćŚÔţĺŘ ďĄÑĂõ SSþŘ ĆŞ ÀþŚśœ ø ťĺŘ ùÀł ùćŚÔţĺŘ ýŘùćÂţÆð ďŚČÆŞ SSþŘ ďć ùÀĂĂîùŘÂÞð ø ŰøŚÔţõ ďŚČÆŞ ücĂãõ øć Ćî ÀĂţÆœŘćüõ Free ýĆÞÜî ěŘ ÂŮĆŞ ,üœŚãõ SSþŘ ôcŚÞŮ üĂãõ SSČóøŘ ĆîöŊ ñŚ§ ,ťĺŘ ùÀł ùćŚÔţĺŘ ćŘěŊ ücĂãõ ĆŞ Free Software ŰcďŚśä ďć Free .ćďŘć ĆĂČõě Ćî üŮďĄň ďć ,ÀłŚŞüõ üœŚźõ üĂãþ öŊ Â cÚþć ücĂãõ ćĄłüõ ďćŚśţõ SSûč ĆŞ ŰďŚśä SSþŘ öcÀþć ŚŞ Ćî øć ěŚŞ Àî Çc śĂŻ ø ćŘěŊcďŘÃêŘôÂœ ćcŚČĂŞ ĆİÂðŘ .ćďŘÀœ ďŘÃêŘôÂœ ľc øÂê ŚŞ üÌìŚĂŮ žČû ćŘěŊcďŘÃêŘôÂœ ŚûöŊ écŘÂÏŘ ťc ČãÞŻ ø ùćĄśœ řČìď Śþ ýěŘĄõ Ýû ŚŞ ĆŻø žČû ĆŞ üóø ,ÀĂłŚŞüõ ÓÜţ£õ ťc þćĄŻĄõ SSþŘ ýĆõŘćŘ ďć ĆİöŊ .ÀœďŘć Řď ćĄą Õ cČìć SS cČœŘĄì ø ÓþďŚãŮ ôŘÀî Âû øć SSþŘ .ćďŘć íŘÂţłŘ Ýû ŚŞ ďŚČÆŞ ďĄÑĂõ ø ,ÀłŚŞüÞœ øć SSþŘ ěŘ ŒþžČû ĆŞ ĆţÆŞŘø ŚňŚÊţąŘ ćÂî ÝČûŘĄą ťś¡ň öŊ ýùďŚŞďć ĆóŚÖõ

.ÀœÂČðüõ ÂÑœ ďć ÂŞďŚî ýŘÂŞ Řď üúŞŚÈõ ëcĄÖ§ ćĄą SS cČœŘĄì ďć Ćî ťĺŘ ýďŘÃêŘôÂœ ýŚûťČóŚãê ôcŚÞŮ

ćŘěŊ ďcŘÃêŘôÂœ À cČóĄŮ À cœøď 4

:ćÂî ÝČÆÖŮ üÜî ýĆţĺć øć ĆŞ öŘĄŮüõ Řď ćŘěŊ ýŚûďŘÃêŘôÂœ

ýŚţĺŘď ďć ø ťîÂł Œþ ÍcĺĄŮ ŚûöŊ ýcÃîÂõ ýĆ£Æœ Ćî üþŚûďŘÃêŘôÂœ :^{14} ùÀłýďŘÁðĆþŚõÂĺ • ćcŘÂêŘ Œc Þî ěŘ ŚûťîÂł SSþŘ Ćî ťÆČœ Âîč ĆŞ üõøÃó .ÀœĄłüõ ÀČóĄŮ ťîÂł öŊ écŘÀûŘ Apache Web Server öŘĄŮüõ Ćţĺć SSþŘ ěŘ .ÀœÂŞüõ öŘøŘÂê ýùÂúŞ ťîÂł öcøÂČŞ ďć ÀĂõĆìĚä

.ćě ñŚůõ Řď PHP ø

ťc þŚÞ§ üţîÂł žČû Íc ĺĄŮ ŚûöŊ ýcÃîÂõ ýĆ£Æœ Ćî üþŚûďŘÃêŘôÂœ :^{15} ýďŘÁðĆþŚõÂĺ öcøÀŞ • ĆãĺĄŮ ÀĂĂîüõ ďŚî öŊ ýøď Ćî ýćŘÂêŘ üc Ê£ł ýĆìĚä ťc úŻ ĆŞ ø ,ćĄłüÞœ ÝČÖţÆõ

.PostgreSQL À cĂœŚõ .ÀĂŞŚþüõ

,ôøć ýĆţĺć ýŚûďŘÃêŘôÂœ :ÝČĺďüõ üúŻĄŮ ÛcŞŚì ýĆţØœ ĆŞ ëĄê ýcÀĂŞĆţĺć ďć ťìć üÞî ŚŞ ŰcďĄň ĆŞ ÂðŘ ø ÀĂŞŚþüõ ĆãĺĄŮ ,ÀłŚŞ ùćŘć ľďŚÔĺ Řď ŚûöŊ üÊ£Èõ ťc îÂł Śþ É£ł ĆîöŊ öcøÀŞ ćŘěŊ ĚõŚî ýcÂŮĄČşõŚî ýŚûÝţÆČĺ ĆŞ üŞŚþťĺć écÀû ĆŞ Řď Śõ ěøďĆŞěøď ,ćĄł ùŚÚœ ŚûöŊ ĆŞ Ûî Œþ SSČĂİ ýøď ÂŞ ďŚî ýŘÂŞ ćŘÂêŘ ýùÃČÚœŘ ćŘěŊcďŘÃêŘôÂœ ø ěŚŞ Àî öcŘďŘÀêÂÏ ěŘ ýďŚČÆŞ .ÀĂĂîüõ ÂŮŒþćÃœ ĆîöŊ ñŚ§ ,ÀĂœŘćüõ üþŚúœ écÀû SS cČĂİ ĆŞ öÀČĺď ďć ťČóĄßÆõ ğc ŚÆ§Ř ø ĆìĚä Řď üþŚûùęøÂŢ SSþŘ ťc ÈŢ ýĆÔÆÜê ĆČŻĄŮ ýŘÂŞ üûŚðÀþć SSČĂİ ĆŞ ýěŚČœ Ćî ÀœŘùćŘć öŚÈœ ýďŚČÆŞ ýŚûťîÂł ,öćĄŞ ğďťĺć ďć öcŚØõŘ Ćî ÀœÀÖţãõ üÌãŞ ,ćÂî ÝČûŘĄą öŚČŞ ÂŮĄÜŻ Ćî ďĄÏöŚÞû .ťÆČœ ŚûďŘÃêŘôÂœ

Eric Raymond^{13}

Funded^{14} Unfunded^{15}

.ťĺŘ ŰďŚźŮ ýŚČœć ďć ťŞŚìď ďć ýŘùÀœÂŞ ïc ÂŞ ,âśĂõ À cî

øŘ .Àûćüõ ôŚźœŘ ýÂÚþć ù cŚðÀþć ěŘ Řď ýÀĂŞÝČÆÖŮ SSþŘ ]2[ćĄą ücŞĚÖœŘ ŠcŚţî ďć ÀœĄÞþď ŒþďŘ :ÀĂîüõ ÝČÆÖŮ Âþě ýĆţĺć øć ĆŞ Řď ćŘěŊ ďcŘÃêŘôÂœ À cČóĄŮ ÍcČ¡õ

ø ,ćĄłüõ ÀČóĄŮ ĆţÆŞ ø ýÃîÂõ Ý cČŮ Œþ ÍcĺĄŮ ýÃîÂõ ďŘÃêŘôÂœ ,ÂţÆŞ SSþŘ ďć :^{16} âõŚŻ ýŚÆČÜî • ěŚŞ öŘÀĂİ ďŘÃêŘôÂœ ýĆãĺĄŮ À cœøď ŰcŚČŐÃŻ ĎĄÞãõ .ćĄłüõ ĆŐŘďŘ öŘÂŞŚî ýĆãõŚŻ ĆŞ ěŚŞ À cî ŚŞ ÂÚþć ěŘ .ćĄłüõ SSČČãŮ ýÃîÂõ Ý cČŮ Íc ĺĄŮ ďŘÃêŘôÂœ ücÜňŘ ücÈõ Íą ,ńĄÊąĆŞ ø ,ťÆČœ ďcŘÃêŘôÂœ ýŘÂŞ ÂÚþć öcŚÆþĄœĆõŚœÂŞ Íc ĺĄŮ Ćî ^{17} üþŚûĆÜňø Ćî ťĺŘ SSþŘ ñÀõ SSþŘ ýŚûĆœŚÈœ ŰcďĄň ďć ø ,ťêŚþ ÀûŘĄ£œ ùŘď ýÃîÂõ âcþěĄŮ ĆŞ üśĺŚĂõ öcŚõě ďć ŚõøÃó ,ćĄł ĆţłĄœ üÜňŘ ýÃîÂõ Ý cČŮ ,ĆňĚą ďcĄÏ ĆŞ .ćĄŞ ÀûŘĄą ÂÆČõ ýÃîÂõ Ý cČŮ Õ cČìć ücĺďŘø ěŘ ÅŢ ŚúĂŮ ,üŞŚþùŘď ćďĄõ SSþŘ ďć) ťĺŘ ÝØ§ ďcøÀň ľŘĆÔČÒø Ćî ÀĂîüõ ÛÞä âõŚŻ ýŚÆČÜî Œþ ýŚÌäŘ À cĂœŚõ

.(ďŘÃêŘôÂœ

Âû .ÀŞŚþüõ ľÂţÆð ěŚŞ ĚõŚî üÎČ¡õ ďć ďŘÃêŘôÂœ ,Ûśì Â cţÆŞ écĚą ÂŞ ,ÂţÆŞ SSþŘ ďć :^{18} ďŘěŚŞ • ÅþĄœĆõŚœÂŞ Âû ø ,ťłŘć ÀûŘĄą ýÂŮÇČŞ čcĄÔœ ,ÀłŚŞ ĆţłŘć ýÂţÈČŞ ücĂê ücþŚœŘĄŮ Ćî Åî ø ĆãĺĄŮ ďć ÂÚþć Åc þĄœĆõŚœÂŞ Âû ýùěŘÀœŘ ĆŞ ,üĂê üc þŚœŘĄŮ ø ÛþŚÞŮ ŰcďĄň ďć ÀœŘĄŮüõ ,ýÃîÂõ ŚţśÆœ üûøÂð ďcŚî ÂŞ ùøĚä ľøď SSþŘ ĎĄÞãõ .ÀłŚŞ ÝČúĺ ŚûďŘÃêŘôÂœ SSþŘ ýcÂČðťúŻ Œc İĄî ø ĆţÆÆð ýŚûŒÞî ÂŞ üĂţśõ ,ÀĂĂî ÂČČçŮ ĚõŚî ťĺŘ SSØÞõ öŚõě ťc łÁð ŚŞ Ćî ĆŞ Řď ýÀþÀŻ öcŚØõŘ ,ěŚČœ Śþ ,ĆìĚä řc śĺ ĆŞ Ćî ťĺŘ ŚČœć Â cĺŚŮÂĺ ěŘ ÓÜţ£õ öcŚÆþĄœĆõŚœÂŞ ĆŐŘďŘ ĆÜňø Œþ řc óŚì ďć ,ýÃîÂõ ýĆ£Æœ ďć řÊœ ýŘÂŞ Řď öŊ ø ,ÀœŘùćÂî ĆêŚőŘ ďŘÃêŘôÂœ ďcŘÃŞŘ ø ^{19} üţÆŢ ťc ĺÂúê ěŘ SSČÚĂĺ ýùćŚÔţĺŘ ÂŞ üØţõ ñÀõ SSþŘ ďć ďŘÃêŘôÂœ ľc ÂţÆð .ÀĂĂîüõ Åî Âû öŊ ďć Ćî ćÂî ĆČśÈŮ ýďŘěŚŞ ĆŞ Řď ÍČ¡õ SSþŘ öŘĄŮüõ ,ĆŞŚÈõ ďcĄÏ ĆŞ .ÀłŚŞüõ ^{20} CVS (ďŘÃêŘôÂœ ćďĄõ SSþŘ ďć) ýÂţÈõ ýŚûěŚČœ âcêď ďć ÀœŘĄŮüõ ,Àûć ĆŐŘďŘ ýÂŮĆŞ ťc ČÔČî ŚŞ ñcĄÊ¡õ

.ćě ñŚůõ Řď Mozilla öŘĄŮüõ Ćţĺć SSþŘ ěŘ .ÀłŚŞ ÝČúĺ Řď ěŚŞ Àî Çc śĂŻ ø ćŘěŊcďŘÃêŘôÂœ ćcŚČĂŞ ù cŚðÀþć SS cČŞ üÜňŘ ëcÂê öŘĄŮüõ ÓþďŚãŮ SSþŘ ĆŞ ĆŻĄŮ ŚŞ ÂŞďŚî ýŘÂŞ ,âśĂõ À cî Â cČČçŮ ø ,ďŚÈţœŘ ,üşî Õ c§ ÐcÔ§ ñcŚśœć ĆŞ ćŘěŊcďŘÃêŘôÂœ ćcŚČĂŞ :ćÂî öŚČŞ ĆœĄðSSþŘ ďŘěŚŞ ñcÀõ öŚÞû üĂãþ ,ďŘÃêŘôÂœ ýĆČúŮ À cœøď öcćĄŞ ěŚŞ ýøď ÂŞ ěŚŞ Àî Çc śĂŻ À cČîŚŮ üóø ,ÀłŚŞüõ .ťÆû ÃČœ ďŘěŚŞ ñcÀõ ěcŚČœÇČŢ ,âśĂõ À cî Â cČČçŮ ø ,ďŚÈţœŘ ,üşî ëcĄÖ§ Ćî ťÆČœ SSţÔð ĆŞ ýěŚČœ .ÀłŚŞüõ

Cathedral^{16}

Patch^{17} Bazaar^{18} Mailing List^{19} Concurrent Versions System^{20}

ďŘÃêŘôÂœ ücĺÀĂúõ 5

Àî ýŚûùęøÂŢ ěŘ ýďŚČÆŞ ďć ,ñŚ§ SSþŘ ŚŞ .ťĺŘ "ĆõŚœÂŞ SS cţłĄœ’ ěŘ ÂŮâČĺø ýŘùěĄ§ ďŘÃêŘôÂœ ücĺÀĂúõ ťÆČœ ôěĎ Ćî ÀþŊüõÂŞ ĆœĄðSSþŘ ü£þďŚŮ ýŚûñŚůõ ěŘ .ćøďüõ öøÂČŞ ø ćĄłüõ ĆţłĄœ ŚúĂŮ ĆõŚœÂŞ ,ěŚŞ ücÜňŘ ýŘÃŻŘ ďŚÊţąŘ ĆŞ Âþě ďć .ćÂČð ďŘÂì âČĺø ýùćŚÔţĺŘ ćcďĄõ ŚŮ ÀłŚŞ ùÀł üĺÀĂúõ ŚÞţ§ ďŘÃêŘôÂœ ĆŞ öŚþŚŢ ďć .ÝČĂîüõ üêÂãõ ěŚŞ Àî ýŚûáŚÞţŻŘ ďć Řď öŊ ÿcþŘď ñcćŚãõ ø ùćÂî öŚČŞ Řď ďŘÃêŘôÂœ ücĺÀĂúõ

.ćÂî ÝČûŘĄą ùďŚłŘ ŚûŰøŚÔŮ SSþŘ ÿcþŚţœ

:ćÂÞłÂŞ ĆœĄðSSþŘ ďŚÊţąŘ ĆŞ öŘĄŮüõ Řď ďŘÃêŘôÂœ ücĺÀĂúõ ýŘÃŻŘ

21 ŚûěŚČœ ćcďøŊÂŞ •

22 ŚûÝţÆČĺ ĳcÎĺ üc§ŘÂÏ •

23 üŐÃŻ üc§ŘÂÏ •

24 ýěŚĺùćŚČŢ •

25 ýÀĂŞĆţÆŞ ø ÕČÔÜŮ •

26 ýŘùěĄ§ ťcÆŮ •

27 üœŚśČţÈŢ •

ýŚûÇ£Ş ĆŞ ĆţÆŞŘø Ç£Ş Âû ,ťĺŘ ÿþŘď ďŘÃêŘôÂœ ücĺÀĂúõ ďć Ćî ďĄÏöŚÞû ,ŘÃŻŘ SSþŘ ćcďĄõ ďć Ç£Ş Œþ ýcěŚĺùćŚČŢ ùŚð ø ÓČňĄŮ SSØČóø ,ÀŞŚþ öŚþŚŢ ŚûöŊ ěŘ ÇČŢ ÀœŘĄŮüÞœ ø ,ťĺŘ ćĄą ěŘ ÇČŢ ďŘÃêŘôÂœ ücĺÀĂúõ À cœøď Â cÚþć {ń \atop c} ŘĄą ěŘ .ćĄł ôŚźœŘ ľěŚČœÇČŢ ýŚûÇ£Ş ôcŚÞŮŘ ěŘ ÇČŢ ťĺŘ SSØÞõ

Marketing Requirements^{21}

System-Level Design^{22}

Detailed Design^{23} Implementation^{24}

Integration^{25} Field Testing^{26}

Support^{27}

.ťĺŘ Û§ŘÂõ ôcŚÞŮ ďć ^{28} üĂČŞěŚŞ ĆŞ ěŚČœ

ŚûěŚČœ ćcďøŊÂŞ 1.5

,ĆõŚœÂŞ ěcŚČœ ćcďĄõ ŰcŚœŚØõŘ ,ÂŞďŚî ýŚûěŚČœ ćcďøŊÂŞ ø ýďøŊćÂð ďŘÃêŘôÂœ ücĺÀĂúõ À cœøď ěŘ ôÀì SSČóøŘ ćcďøŊÂ Ş À cĂ ĺ Ć î ,ćĄ łü õ Ć ţ łĄ œ ýćŚ þě Ý cź § ĎĄ Þ ã õ Ć Ü §Â õ SS þŘ ďć .ť ĺŘ ř ÏŚ £ õ SS cţ êŚ þ ø

.ťĺŘ Ćţĺć öŊ ěŘ ^{29} (MRD)ŚûěŚČœ

ôŚźœŘ ^{30} ýÂśą ù cøÂð Śþ üţÆŢ ťc ĺÂúê Õ cþÂÏ ěŘ ĎĄÞãõ ěŚŞ Àî ýŚûùęøÂŢ ďć MRD ýĆČúŮ ,ÀĂţĺÂêüõ ôŚČŢ ôŘÀõ ťÈðÂŞ ø ťêď ŰcďĄň ĆŞ ďŘÃêŘôÂœ öcŚðÀĂûćĆãĺĄŮ ø öŘÂŞďŚî öŊ ďć Ćî ,ćĄłüõ öÀČĺÂœ áŚÞŻŘ ĆŞ ŰŚìøŘ ěŘ ýďŚČÆŞ Ćî ťĺŘ öŊ ñÀõ SSþŘ řcþŚãõ ěŘ .ÀĂĺÂŞ áŚÞŻŘ ĆŞ ÀĂĂîüõ üãĺ ø ĆţêÂð ďŚî SSþŘ ýĄÜŻ ćĄłüÞœ üãĺ üţ§ ćďŘĄõ ěŘ ýďŚČÆŞ ďć .ćĄŞ ÀûŘĄą "^{31} Àî Ý cČÆÖŮ’ ücĂãõ ĆŞ

.ťĺŘ SSØÞõ Â cČè Ûî ĆŞ Śþ ,ćĄł

ÝţÆČĺ ĳcÎĺ üc§ŘÂÏ 2.5

ěŘ éÀû .ÀþŊüõ ťĺć ĆŞ Ýû ŚŞ ŚûöŊ ÍcŞŘøď ø "ŚûĆœŚÞČŢ’ ğc ŚĺŘ ÂŞ ÝţÆČĺ Óc ČňĄŮ ĆÜ§Âõ SSþŘ ďć ýŘÂŞ ôěĎ ďcŚî SS cČÞ£Ů Åşĺ ø ,đÂÏ öcćĄŞ ŘÂŻŘ ÛcŞŚì ø üţĺďć ěŘ öŚĂČÞÏŘ ñcĄÊ§ ŘÀţŞŘ ĆÜ§Âõ SSþŘ

.ťĺŘ öŊ ýcěŚĺùćŚČŢ

ŰcďĄň ĆŞ ÝţÆČĺ ø ,ćďŘÀœ ćĄŻø ĆÜ§Âõ SSþŘ ýďŘÁðĆþŚõÂĺ öcøÀŞcěŚŞ Àî ýŚûùęøÂŢ ďć ĎĄÞãõ ôcÀä âìŘĄõ ďć .(ďŘÃêŘôÂœ ćcĄą À cĂœŚõ) ÀĂîüõ ÂČČçŮ öŚõě ťc łÁð ŚŞ đÂÏ Śþ ø ććÂðüõ ü§ŘÂÏ üĂÞő ćcĄŻø ôcÀä üóø ,ćÂî üÔ£õ ýÀ§ ŚŮ ùÂśą öcŚÆþĄœĆõŚœÂŞ ďcĄÌ§ ŚŞ öŘĄţŞ ÀþŚł Řď QA Śþ MRD ćcĄŻø

.Àł ÀûŘĄą ćøÀ¡õ ùęøÂŢ ťcČÔČî ĚÞä ,ÝţÆČĺ üc§ŘÂÏ

üŐÃŻ üc§ŘÂÏ 3.5

ĆŞ ÂÚþÀØþ ĆŞ ŚûĆœŚÞČŢ ücÚţÆŞŘø ýŚûĆÎŞŘď ø ,öŊ ÍcĺŘø ĆÞÜŻ ěŘ ,ĆœŚÞČŢ Âû ŰcŚČŐÃŻ ĆÜ§Âõ SSþŘ ďć ÀûŘĄą ĆŐŘďŘ ÃČœ GANT ø PERT ýŚûďŘćĄÞœ ĎĄÞãõ ĆÜ§Âõ SSþŘ ücŻøÂą öcŘĄĂä ĆŞ .ÀþŊüõ ťĺć ü ŐÃ Ż üc §ŘÂ Ï À cĂ ĺ Řď ěŚ ê SS þŘ üc ŻøÂ ą .À Ă ûćü õ öŚ È œ Řď Ś ûďŚ î ôcŚ ź œŘ řc Č ŮÂ Ů ø öŚ õě Ć î ,À ł

Review^{28} Marketing Requirements Document^{29}

Newsgroup^{30} Code Splits^{31}

.ÀĂõŚœüõ ^{32} (DDD)

SSþŘ ôćÂõ ěŘ ýďŚČÆŞ .ťĺŘ üŐÃŻ üc§ŘÂÏ ,ĳþÂÔŮ ĆŞ ěŚČœ ø ďŘÁðĆþŚõÂĺ SS cţłŘÀœ ÿcþŚţœ ÂÚþć ěŘ Œþ ĆŞ Ćî ÝœŘćüõ’ :ćÂČðüõ ÂČŹŚŮ ýěŚĺùćŚČŢ ěŘ ü§ŘÂÏ ÛÞä ďć ĎĄÞãõ ŚõŘ ,ÀœďŘć ťĺøć Řď ĆÜ§Âõ öŊ ěŘ ÀþŚŞ ĆţśóŘ Ćî ,ťÆČœ ýďŚśŻŘ ÃČœ ŚûÍĺŘø ýcěŚĺÀĂţÆõ ".ÝÆþĄœüõ üØþ ÅŢ ,ôďŘć ěŚČœ ÂĺďŚŢ ŚŞ üÏŚśŮďŘ žČû ø ÀłŚŞ ùÀł ùćŚÔţĺŘ üÜąŘć ýùćŚÔţĺŘ ýŘÂŞ ÍÖê ýŘĆœŚÞČŢ ÂðŘ üţ§ .ÀČÈî ťóŚźą À cî ŰďĄň Âû ďć Ćî ŘÂİ ,ÀœĄł ýěŚĺÀĂţÆõ öŊ ýŚûÍĺŘø ÀþŚŞ Ýû ěŚŞ ,ÀłŚŞ ĆţłŘÀœ ŽďŚą ýŚČœć öŘÂÚþć Íc ĺĄŮ ćÀźõ ýùćŚÔţĺŘ ñcŚÞţ§Ř ,ÀłŚŞ ÂţÈČŞ ŰŘÀĂţÆõ Ćİ Âû ø ťĺŘ ôćÂõ ďcŚČţąŘ ďć âśĂõ

.ćĄłüõ ÂŮÇČŞ

ýěŚĺùćŚČŢ 4.5

Œc İĄî ÛcÞä öŚÞû SSþŘ .ÀœĄłüõ ýěŚĺùćŚČŢ Ç£Ş SSþŘ ďć DDD ďć ùÀł đÂÎõ ýŚûĆœŚÞČŢ ćĄłüõ ùÀþć ùŚð ĆœŚÔĺŚţõ .ťĺŘ ďŘÃêŘôÂœ ücĺÀĂúõ đcøď ø řÜì Ćî ,ÀłŚŞüõ ĆœŚÞČŢ Âû ücÆþĄœĆõŚœÂŞ ñŚ§ SS cČä ďć Ćî üóŚ§ ďć ,(ćĄłüõ ĆţêÂð ćŚþ) ćĄłüõ ÅþďÀŮ Ćî ťĺŘ üÈ£Ş ŚúĂŮ Ç£Ş SSþŘ Ćî

.ćěĄõŚČŞ ćĄą ĆŞ üţ§Řď ĆŞ ÀœŘĄŮüõ É£ł Ćî ťĺŘ ďŘÃêŘôÂœ ücĺÀĂúõ ěŘ üÈ£Ş ŚúĂŮ

ťĺøć ĆÞû ěŘ ÇČŞ öŚÆþĄœĆõŚœÂŞ Ćî ťĺŘ üţÞÆì ýěŚĺùćŚČŢ .ťĺŘ ďŚî Çc £ŞŰÁó ťc ÞÆì SSþŘ SS cţłĄœ öcŚØõŘ .ÀĂłŚŞ ŠŘĄą ÀĂœŘĄŮüõ Ćî üõŚÚĂû ,ćďŘćüõ ùŚÚœ ďŘÀČŞ Řď ŚûöŊ ĆİöŊ ťĺŘ SSþŘ .ÀœďŘć É£ł ĎĄÞãõ .ťĺŘ ùćĄŞ ěŚŞ Àî ďcŘÃêŘôÂœ ľc ÂţÆð ýŚûľĚŮ ôcŚÞŮ ýŘÂŞ ĆČóøŘ ýŚûùÃČÚœŘ ěŘ üØþ Àî {ŋ \atop c} Âãõ ĆŞ ĆĂČõě SSþŘ ďć Řď ćĄą ýŚûćďøŊťĺć SSþÂŮÀþÀŻ ÀœŘĄŮüõ ěŚŞ Àî ýùęøÂŢ ďć Àî SS cţłĄœ ŚŞ

.ćďŘÁÚŞ öŘÂÚþć À cþć

ýÀĂŞĆţÆŞ ø ÕČÔÜŮ 5.5

ŚŞ ø ÀœĄłüõ ĆêŚőŘ À§Řø À cî ťc ąďć Œþ ĆŞ ø ĆţÈð âÞŻ ùÀłýěŚĺùćŚČŢ ýŚûĆœŚÞČŢ ĆÜ§Âõ SSþŘ ďć ø ,ÀœĄłüõ ĆţłĄœ ĆœŚÞČŢ Âû ýŘÂŞ řÊœ ýŚûďŘÃêŘôÂœ ôøÃó ŰcďĄň ďć .ÀœćÂðüõ ŒĂČó ø ÛþŚşõŚî Ýû

.ÀœĄłüõ ťÆŮ ďŚÑţœŘ ćcďĄõ ýŚûďŚţêď ø ŚûĆœŚÞČŢ ĆŞ ćøďø ŕcŚÖœ ôcŚÞŮ ,ŒČŮŚõĄŮŘ ťcÆŮ Ý cţÆČĺ ÍcĺĄŮ ěŘ öŚĂČÞÏŘ ñcĄÊ§ ,ŚÞĂûŘď ýĆ¡Ôň ÀĂİ SS cţłĄœ ÛcõŚł ĎĄÞãõ ĆÜ§Âõ SSþŘ ěŚŞ Àî ýùęøÂŢ Œþ ďć Œþ SS cţłĄœ ,Makefile ýcďŚîÃČÞŮ ,ğďťĺć ďć ýŚûùŚðťĺć Óc Üţ£õ ácŘĄœŘ ýøď ĳČ¡ň ÛcþŚşõŚî ŚţþŚúœ ø ,FTP ťc þŚĺ Œþ ýøď öŊ öcćŘć ďŘÂì ,^{33} Tarball ýùćÂÈê ýĆţÆŞ SS cţąŚĺ ,README ťc ÆŮ ÀþÀł ĆŻĄţõ Ćî ĆœĄðöŚÞû .ÀłŚŞüõ ýÂśą ýŚûùøÂð ø üţÆŢ ýŚûťĺÂúê ďć üõĄÞä öcĚäŘ ćĄŻø ÃČœ ŚĂůţĺŘ) ťĺŘ Œśĺ ďŚČÆŞ ěŚŞ Àî ýŚûùęøÂŢ ďć ĎĄÞãõ Ç£Ş SSþŘ .ťÆČœ ďŚî ďć ÝţÆČĺ öŘÀĂİ ĆÜßÆõ SSþŘ ,Àł ÀĂûŘĄą öŚČŞ Ćî ýÂÚþć ÛcþĎć ĆŞ ĆţśóŘ .(PostgreSQL ø Perl À cĂœŚõ ,ćďŘć

Detailed Design Document^{32}

.ÀþŊüÞœ ŠŚÆ§ ĆŞ üðďÃŞ Óc ãő

ýŘùěĄ§ ťcÆŮ 6.5

ťĺŘ ôěĎ üŮÀõ ěŘ ÅŢ .ÀĂĂîüõ ñĄÊ¡õ ěŘ ùćŚÔţĺŘ ĆŞ áøÂł ćĄą öŘÀĂõďŚî ŘÀţŞŘ ĆÜ§Âõ SSþŘ ďć Ý cţÆČĺ ĆŞ ťśÆœ üŮøŚÔţõ À cþć ŚûöŊ ŘÂþě ,ÀĂûć ôŚźœŘ Řď ďŚî SSþŘ ÃČœ ùęøÂŢ Ý cČŮ ěŘ ŽďŚą ýćŘÂêŘ Ćî

.Àł ÀĂûŘĄ£œ ÓÈî ÝČŮ ÍcĺĄŮ Ćî ćÂî ÀĂûŘĄą ÓÈî Řď üŮĎŚØłŘ ø ÀœďŘć ùÀłýěŚĺùćŚČŢ

ÛcČóć .ÀœďŘć ŚČœć ďć Řď ýŘùěĄ§ ťc ÆŮ Ý cţÆČĺ SSþÂţúŞ ýďŘÁðĆþŚõÂĺ öcøÀŞcěŚŞ Àî ýŚûďŘÃêŘôÂœ ćcďĄõ ďć ÂŮĆœŚţĺøć ďŚČÆŞ ÀĂłŚŞ ùćŘÀœ ñĄŢ ďŘÃêŘôÂœ ýŘÂŞ üţìø öŘÂŞďŚî Ćî ťĺŘ SSþŘ öŊ ýùćŚĺ ďŚČÆŞ ø ,ÀœěŚĺ éÂÏÂŞ Řď řČä ćĄą ÀĂĂîüõ üãĺ öŚØõŘ À c§ ŚŮ üţ§ ø ,ÀĂĂîüõ ćďĄąÂŞ ďŘÃêŘôÂœ ŰcĚØÈõ

.ÀĂţĺÂÔŞ Řď ĆÏĄŞÂõ ýĆÜňø

Ćî ,ťĺŘ ÂÚþć öcŚÆþĄœĆõŚœÂŞ Ý cÈİ ťÔŻ ŚûÀň ÍcĺĄŮ Àî ücĂČŞěŚŞ ěŚŞ Àî ýŚûùęøÂŢ Â cÚþć À cþŘĄê ěŘ ñĄŢ ćŘÂêŘ SSþŘ ěŘ üÌãŞ .öŊ ýŘÂŻŘ Õ cþÂÏ ěŘ Ćœ ,ÀĂŞŚþüõ Řď öŊ ŰcĎŚØłŘ ĆõŚœÂŞ À cî öcÀœŘĄą ù cŘď ěŘ ùÀłŘÀČŢ ŰcĎŚØłŘ ćďŘĄõ řcÜèŘ ďć .ÀĂĂî üĺďÂŞ ,Àî öcÀœŘĄą ŚŞ ,ťČĂõŘ Â cÑœ ěŘ Řď ýďŘÃêŘôÂœ ŚŮ ÀœÂČðüõ

.ÀœĄłüõ ľďŘÃð ùÀĂÆþĄœ Ý cČŮ ĆŞ ÕþÂÏ SSþŘ ěŘ

üœŚśČţÈŢ 7.5

ýùćŘć ù cŚÚþŚŢ Œþ ďć ÀþŚŞ ÀœĄłüõ ťêŚþ öŊ ěŘ ÅŢ ø ýŘùěĄ§ ťc ÆŮ Ûc§ŘÂõ ďć ÅŢ Ćî üŮĎŚØłŘ ,ćĄł ďŘÁðŘø ĆÏĄŞÂõ SS cČĺÀĂúõ ěŘ üØþ ĆŞ ñŚØłŘ Âû ÀþŚŞ öŊ ěŘ ÅŢ .ÀœĄł ýďŘćĆÚœ ø Íśő ýÃîÂõ ùÀł ťśŹ ùćŘć ù cŚÚþŚŢ ďć ÛØÈõ ø Û§ ù cŘď SSþŘ .ÀĂî ĆŐŘďŘ ÛØÈõ âcêď ýŘÂŞ üÜ§ ù cŘď ĆÏĄŞÂõ ğc ÀĂúõ ø ĆŞ ĆÏĄŞÂõ ñcŚØłŘ ücþŚĺŚĂł ďć Ćî ÀþÀŻ ^{34} öcĄõěŊ ýùćŘć ĆœĄðÂû SSČĂżÞû ,ÀœĄłüõ ýďŘćĆÚœ ø ùÀĂþŊ ďć ÛØÈõ SSþŘ ćcÀźõ ěcøÂŞ ěŘ ŚŮ ććÂðüõ ĆêŚőŘ öĄõěŊ ýŚûùćŘć ýĆäĄÞźõ ĆŞ ,ćøďüõ ďŚî

.ćĄł ýÂČðĄÜŻ

Ćî ÀĂĂîüõ ÛÞä ýŘĆêÂ§ öŚĂİöŊ ÃČœ ýďŘÁðĆþŚõÂĺ öcøÀŞcěŚŞ Àî ýŚûùęøÂŢ ěŘ ýďŚČÆŞ ùěøÂõŘ üŞŘĄŻ ,ÂŮŒİĄî ýŚûùęøÂŢ řc ÜèŘ ďć üóø ,PostgreSQL À cĂœŚõ ,ÀœďŘÀœ üœŚśČţÈŢ ďć üÊÖœ žČû SSþŘ ÂÚþć écÂÏ ěŘ .ťĺŘ ".ÀČÈ£śŞ ,ùøŘ’ :À§ ďć ,ťêÂð ÀČûŘĄą ñŚØłŘ Œþ ľc ďŘÃð ÛcŞŚÖõ ďć Ćî

.tar.gz^{33} Test Data^{34}

.ÀĂîüõ ěŚŞ ÛÖţÆõ ďcŘÃêŘôÂœ ücœŚśČţÈŢ ýŚûÝČŮ ýŘÂŞ Řď ýŚŻ üœŚśČţÈŢ SS cţłŘÀœ

ěŚŞ Àî ďcŘÃêŘôÂœ ÀĂİ ÂŞ üûŚÚœ 6

ŚŞ .ćÂî ÝČûŘĄą ùďŚłŘ ěŚŞ Àî Çc śĂŻ ¿cþďŚŮ ďć üœÀÈœ ľĄõŘÂê ø řóŚŻ ýùÀœøÂŢ ÀĂİ ĆŞ Ç£Ş SSþŘ ďć ťĺć ěŚŞ Àî ýcďŚźŮ ýŚûĆśĂŻ ćcďĄõ ďć üœŘøŘÂê řc óŚÎõ ĆŞ öŘĄŮüõ ,ŚûťČãìŘø SSþŘ ĆŞ ÕČìć üûŚÚœ

.ťêŚþ

Linux 1.6

ø ü§ŘÂÏ Ćî ĆţÆû SSþŘ .ťÆœŘć ěŚŞ Àî Çc śĂŻ ťc ČÖêĄõ SSþÂŮïďÃŞ öŘĄŮüõ Řď ÅØĂ cČó ýĆţÆû öcŚÆþĄœĆõŚœÂŞ ŠcÁŻ ŚŞ ťÆœŘĄŮ üÞî ŰcÀõ ěŘ ÅŢ ,Àł áøÂł ^{35} ěÀóŘøÂŮeÅĂ cČó ÍĺĄŮ öŊ ýcěŚĺùćŚČŢ ñcÀõ ďć ÅØĂ cČó .ććÂð ÛþÀśŮ ÀĂõŰďÀì ďŚČÆŞ ÛcõŚä Ý cţÆČĺ ýĆţÆû Œþ ĆŞ ŚČœć Â cĺŘÂĺ ěŘ ýÀĂõŰďÀì ĆŞ ýÂČČçŮ Âû ücþŚúœ À cČþŚŮ ,öŊ ťc ĺďć ýcďŚÞãõ ø üţąŘĄœŒþ ÐÔ§ ýŘÂŞ üóø ťĺŘ ùÀł ÀČóĄŮ ďŘěŚŞ

.ťĺŘ ěÀóŘøÂŮe{É \atop c} £ł ýùÀúä

KDE ø QT 2.6

ýĆþŚŢ ÂŞ Ćî ćĄŞ SSþŘ öŊ ÛcØÈõ ŚõŘ ,Àł đÂÎõ ÅØĂČó ücØČêŘÂð ÍcĺŘø SSČóøŘ öŘĄĂä ĆŞ KDE Ý cţÆČĺ Àî Çc śĂŻ ŚŻSSþŘ ďć .ćĄśœ ěŚŞ Àî ÛõŚî ďcĄÏ ĆŞ öŊ ěcĄźõ Ćî ćĄŞ ùÀł ĆţłĄœ QT ôcŚœ ĆŞ ýŘĆœŚąŠŚţî Â cČČçŮ ĆŞ ďŘćŘø Řď ťîÂł öŊ ćďøŊ ćďŘø TrollTech üĂãþ QT ýùÀĂĂî ÀČóĄŮ ťc îÂł ĆŞ Ćî ýďŚÈê ŚŞ ěŚŞ GNOME ýùęøÂŢ øć ácøÂł ÃŻ ćĄśœ ýÃČİ ŚûďŚÈê SSþŘ .ćÂî ěŚŞ Àî ĚõŚî ěcĄźõ Œþ ĆŞ QT ěcĄźõ ĆţłĄœ ěŚŞ Àî ŚŞ ŘÀţŞŘ ěŘ Ćî ćĄŞ KDE ñcćŚãõ üØČêŘÂð Íc ĺŘø Œþ GNOME Ý cţÆČĺ .Harmony ø ŰcďĄň ďć Ćî ,ćĄŞ QT ýĆœŚąŠŚţî Íc ĺŘø ěŘ ćŘěŊ ø ÛÖţÆõ ýcěŚĺùćŚČŢ Œþ Harmony ø ,Àłüõ øć SSþŘ ĆîSSþŘ ěŘ Ûśì ďŚČÆŞ TrollTech ťc îÂł .ÀĂî éÁ§ QT ĆŞ Řď KDE ücÚţÆŞŘø ťÆœŘĄŮüõ ôŚÞŮŘ

.Àł ÝČÜÆŮ ÀĂĺÂŞ ùćŚÔţĺŘ ÛcŞŚì ýĆÜ§Âõ ĆŞ ùęøÂŢ

Linus Torvalds^{35}

Netscape 3.6

ýĆźČţœ Ćî ,ćŘć öŚÈœ ĆìĚä ćĄą À cî öcćÂî ěŚŞ ĆŞ ,]2[ÀœĄÞþď ŒþďŘ ŰcĎŚÖõ ýĆãóŚÎõ ŚŞ ťîÂł SSþŘ ťĺŘ ÅþĄœĆõŚœÂŞ öŘďŘÃû ýcďŚîÝû ÛcňŚ§ ÂðďøÂõ SSþŘ .ťĺŘ Mozilla ýŚœŘĄŮ ďŚČÆŞ Â cðďøÂõ ěøÂõŘ öŊ

.ťĺŘ ùÀł ÀČóĄŮ ďŘěŚŞ ñcÀõ ďć ĚõŚî Ćî

RedHat 4.6

ĆŞ ťĺŘ ĆţÆœŘĄŮ ěŚŞ Àî ÂŞ üĂţśõ ýcďŚźŮ đÂÏ ŚŞ Ćî ùćĄŞ üþŚûťîÂł SSþÂŮÕêĄõ ěŘ üØþ ťîÂł SSþŘ ğc ŚĺŘ ÂŞ Řď ćĄą ťcîÂł ýŚĂśõ RedHat ÛcõŚäÂþÀõ ø ÅĺĄõ ,^{36} ÙœŚþ ŠŚŞ .ÀĺÂŞ ŚØþÂõŊ ğc ďĄŞ ďcŘěŚŞ Ćî ťêÂð ŚŻSSþŘ ěŘ Řď ćĄą ýùÀþŘ øŘ .ćŚúœ ÅØĂČó ýĆĂČõě ďć ùćŚţêŘŚŻ ø ôŚœľĄą íc ďŚõ Œþ ýĆŐŘďŘ ěŘ üţÔÚĂû ćcĄĺ Ćî ÀĂţÆû üþŚûťîÂł üóø ,ťĺŘ ğďťĺć ďć ø öŘěďŘ ďŚČÆŞ üœÀČõŚłŊ ŠcŊ ĆîöŊ ŚŞ ,ýěŚĺÛČśõĄŮŘ ýŚûťîÂł ćcďĄõ ďć ďĄÏSSČÞû .ÀœďøŊüõ ťĺć ĆŞ üœÀČõŚłŊ ŠcŊ ýĆŐŘďŘ ø ýÀĂŞĆţÆŞ ďć ñĄŢ öŚłďŚśţäŘ ø íďŚõ ù cŘď ěŘ ø ,ĆţÆŞ Ýû ĆŞ Řď ŰŚãÎì ĆØÜŞ ,ÀœěŚĺüÞœ Řď öŚłŰŚãÎì ćĄą Ćî ø ěŚŞ Àî ďcŘÃêŘôÂœ ÀňďŚúİ ěŘ ÇČŞ ýcÀĂŞĆţÆŞ ø ýďøŊâÞŻ ĆŞ áøÂł ÃČœ RedHat ÅŢ .ÀœďøŊüõ

.ÀĂî řÆî ýćŚþě ýĆþŚõÂĺ ø ďŚśţäŘ ťÆœŘĄŮ ùŘď SSþŘ ěŘ ø ćÂî À§Řø üõŚœ ťc¡Ů ŚûöŊ ýĆŐŘďŘ

Apache Foundation 5.6

Sun ťc îÂ ł ěŘ Řď ćĄ ą üc óŚ õ üc œŚ ś Č ţ È Ţ ,ćěŘćÂ Ţü õ ćŘěŊ ýŚ ûďŘÃ êŘôÂ œ À cČ óĄ Ů Ć Ş Ć î ùøÂ ð SS þŘ .ÀĂîüõ ÀČóĄŮ ŘøŚŻ öcŚŞě ĆŞ ýÀĂõŰďÀì ďŚČÆŞ ýŚûďŘÃêŘôÂœ öŊ ýŘěŘ ďć ø ćÂČðüõ Microsystems

.ćďøŊüõ ťĺć ĆŞ ŘøŚŻ öcŚŞě ĆŞ öŚÆþĄœĆõŚœÂŞ ýĆìĚä Çc þŘÃêŘ ěŘ Řď ćĄą ćcĄĺ ÃČœ Sun ťcîÂł

PostgreSQL ø MySQL 6.6

ñcÀõ ďć MySQL Ćî ŰøŚÔŮ SSþŘ ŚŞ ,ÀĂłŚŞüõ ěŚŞ Àî ýcŘĆÎŞŘď ýùćŘć ýŚûùŚÚþŚŢ øć Âû ÝţÆČĺ øć SSþŘ ø ěŚŞ ĚõŚî ŰcďĄň ĆŞ PostgreSQL üóø ,ćĄłüõ ÀČóĄŮ ÃîÂÞţõ Ý cČŮ Œþ Õ cČìć ŰcďŚÑœ ťc ¡Ů ø ŚÆČÜî À cî ťc ČÔČî Řď ćĄą ýĆãĺĄŮ ñcÀõ ÛcňŚ§ MySQL ĆîöŊ ŚŞ Ćî ťĺŘ řóŚŻ .ďŚČÆŞ öcŚÆþĄœĆõŚœÂŞ ÍcĺĄŮ üœŚśČţÈŢ Řď Transaction ĆÜÞŻ ěŘ ,ùćŘć ù cŚÚþŚŢ Œþ ýĆþŚŢ ŰcŚœŚØõŘ ěĄĂû ÛÞä ďć üóø ,ÀœŘćüõ ÂŮĎŚŞ ýŚûťČÜŞŚì ýŘďŘć ,ťĺŘ ľÂţÈð ĆŞ øď ýćŚþě ďŚČÆŞ ťcäÂĺ ŚŞ PostgreSQL ,ÂÚþć ýĄĺ ěŘ .ÀĂîüÞœ

Bob Young^{36}

.ÀłŚŞüõ Oracle À cĂœŚõ üţÞČìöŘÂð ýŚûÝţÆČĺ ŚŞ ĆÆþŚÖõ ÛcŞŚì öĄĂîŘÝû ø ,ÀłŚŞüõ ýďŚČÆŞ

ĆźČţœ

ŰcŚìøŘ ďć ø ĳþÂÔŮ ťc úŻ ŚúĂŮ ěŚŞ Àî ñcÀõ ďć ďŘÃêŘôÂœ ýĆãĺĄŮ Ćî ćďŘć ćĄŻø üêŚî À cûŘĄł ùěøÂõŘ ěŘ ýďŚČÆŞ ýŘÂŞ ýÀŻ ýÀþÀ¡Ů öĄĂîŘÝû ďŘÃêŘôÂœ À cČóĄŮ ľc øď SSþŘ .Àł ÀûŘĄ£œ ôŚźœŘ ťèŘÂê ñĄŢ öŚłíďŚõ ø ďŚśţäŘ ýøď ěŘ ěĄĂû ,âþÀŞ ýŚûĆÈþÀœŘ ćcŚźþŘ Ý cź§ ěŘ ÇûŚî ŚŞ Ćî ÀłŚŞüõ ŚûťîÂł Řď üêŚî öcŚõě ÇśĂŻ SSþŘ Ćî ŘćÂê ýŚČœć ďć ø ,ťĺŘ öŘĄŻ ďŚČÆŞ ěĄĂû ěŚŞ Àî Çc śĂŻ .ÀœďøŊüõ ďć ,ÀĂî ĆőÂä ďŘěŚŞ ĆŞ Řď ćĄą {É \atop c} ÖœüŞ ŰcĎĄÊ¡õ ø ťĺŘ ĆţêŚþ ÓÜţ£õ ýcÀČÜî ýŚûÝţÆČĺ À cČóĄŮ ýŘÂŞ ýďŘÃêŘôÂœ ýŚûťîÂł řc óŚè ýŘÂŞ üÚţÆØłďø ŚŞ ĆÜŞŚÖõ ýŘÂŞ ñÀõ SSþŘ ĆŞ öćďøŊ ýøď ÃŻ üûŘď .ÀþŚÞœüõ ýďøÂő ďŚČÆŞ ñÀõ SSþŘ ďć ďŘÃêŘôÂœ ücĺÀĂúõ ýŚûľøď ýĆĂČõě ďć ĆãóŚÎõ ŘÁó .ÀœŚõüÞœ

âŞŚĂõ

[1] DiBona, Chris et al, Open Sources, Voices from the Open Source Revolution, 1st ed, 1999, O’Reilly.

[2] Raymond, Eric S., The Cathedral & The Bazaar, Revised, 2001, O’Reilly.

[3] Stallman, Richard, Free as in Freedom, 1st ed, 2002, Sam Williams.

[4] Torvalds, Linus and David Diamond, Just for Fun, 1st ed, 2001, HarperBusiness.